[Hallc_running] [Revised Logentry] RC Daily Update
sjjooste at jlab.org
sjjooste at jlab.org
Mon Oct 15 14:20:03 EDT 2018
- by lassiter on Mon, 10/15/2018 - 13:52.
Clarification on HMS (not) quenching - The HMS Dipole and Q3 fast dump of current was most likely due to the Hall A issues, resulting in a back pressure on the current lead coolant mass flow lines. This generates a fast discharge interlock. No quenching of the HMS magnets has ever occurred to date.
Once accelerator realized they couldn’t deliver 37uA, things ran mostly stable at 35uA during swing and owl. Today during day we won’t have physics beam due to planned beam studies aimed at fixing the issues to deliver high current to Hall C. You can find the beam current history for the last 24 hours here: https://logbooks.jlab.org/entry/3609814
More info about the beam tuning operations here: https://logbooks.jlab.org/entry/3609479.
Yesterday around 7:50pm, the HMS Dipole and Q3 suddenly lost current. Everything was restored back to normal at 9:10pm, so we lost 80min. The shift crew noticed peaks on the warm return when the magnets lost current https://logbooks.jlab.org/entry/3609532. I learned this morning that this coincided with the moment where Hall A warmed up their target, which is most likely what caused the HMS magnet issues.
UPDATE: Steve Lassiter clarified that the fast discharge interlock functioned properly and was most likely triggered by the Hall A issues.
This morning during the beam studies, the SHMS DCs started to trip around 60uA (as expected. We switched of the SHMS DC HV for now. We also masked the alarms in the alarm handler. https://logbooks.jlab.org/entry/3609751
It looks like the JT-4 valve for the HMS Dipole LN2 level is responding again today, after 5 days of being stuck. I talked to Andy and we decided to go in for controlled access tomorrow during the RF recovery to replace the motor to prevent this from happening in the future. Andy expects this to take about 5-10 minutes.
As of right now, we have collected 23% of the statistics goal for LH2 on Q2=5.5, W=3.02, theta_Shms=6.49 during 3 shifts.
Tonight during swing we will most likely interrupt running this setting to conduct the luminosity scan. Tentative run plan for this scan can be found here: https://hallcweb.jlab.org/wiki/index.php/Run_Plan_for_Kaon_LT_Luminosity_Scan
We will then resume taking production data on the same kinematic point for another 2-3 shifts , or until we reach 50% of the statistics goal.
Note that this time allocation excludes the time needed for todays beam studies, the luminosity scan and tomorrow’s RF recovery.
More info about the Kaon LT status here: https://hallcweb.jlab.org/wiki/index.php/Kaon_LT_Status_Page
There will be an RC meeting today at 4:30 pm on the second floor of the counting house.
After the meeting you can find the talking points here:
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-18.104.22.168
Meeting ID : 861439920
To join via phone :
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3609815
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running